OpenStudy (anonymous):

Find the missing coordinate to complete the ordered pair. Equation : 3x + 2y = -2 (-1, )

OpenStudy (goformit100):

Use Substitution Method for solving this equation.

OpenStudy (anonymous):

How?

OpenStudy (goformit100):

We know the value of x coordinate. Do you agree ?

OpenStudy (anonymous):

Yes. 3(-1) + 2y = -2 -3 + 2y = -2 And then? How could i determine the value of y coordinate?

OpenStudy (goformit100):

-3 + 2y = -2 add +3 on both sides. What you get ?

OpenStudy (goformit100):

-3 + 2y + 3 = -2 +3 2y =1 Now divide both side by 2 we get y = 1/2 ................... Ans. Hence required coordinate is (-1, 1/2)

OpenStudy (goformit100):

Did you get the process ??

OpenStudy (anonymous):

Yes :D Thank you very much!

OpenStudy (goformit100):

:) you're welcome